# Verbosity-Aware Rationale Reduction: Effective Reduction of Redundant Rationale via Principled Criteria

### v2

@@ Appendix A Additional Datasets Analysis
As shown in Section 3, the Negative Log-Likelihood (NLL) tends to rise with increasing size of ${r_{i}}$ in random and back configurations, compared to the front configuration, as illustrated in Figure 7 this trend is consistent across both the GSM8K and StrategyQA datasets.
@@ Appendix B VARR Implementation
Appendix C Experimental Details
For all experiments, we employ the AdamW optimizer Loshchilov and Hutter (2019), configured with a weight decay of 0.005. For the Mistral 7B model, we utilize an effective batch size of 12 with gradient accumulation set to 3, while the smaller Llama3.2 models use an effective batch size of 15. A constant learning rate of $5\times 10^{-6}$ is applied across all datasets, with bfloat16 precision. For Multiple-Choice and True/False tasks, complete sets of non-correct labels are employed to configure in-batch negatives to enhance the stability of the verbosity evaluation. To ensure a fair comparison, all baselines and methods are executed on a single A100 GPU with 80GB of memory for up to 5 epochs or 36 hours, whichever is reached first.
@@ Appendix D Details and Statistics of Datasets
GSM8K (Cobbe et al., 2021a) Arithmetic Cobbe et al. (2021b) Number 6999 472 1319
@@
StrategyQA (Geva et al., 2021) Commonsense Sileo (2024) T/F 832 114 344
Table 4: Comprehensive statistics of the datasets used in our experiments are provided. GSM8K and MathQA are sourced directly from their original datasets, while the remaining datasets were obtained from Kim et al. (2023) and Sileo (2024) to access complete reasoning paths.
For our experimental analysis, we carefully selected a diverse set of five datasets used in prior works (Deng et al., 2023, 2024a; Liu et al., 2024; Yu et al., 2024b; Yin et al., 2024). To ensure explicit sentence boundaries, all datasets were processed using gpt4o-mini (OpenAI, 2024) to establish clear sentence demarcation (e.g., ‘He bikes 20*2=<<20*2=40>>40 miles each day for work So he bikes 40*5=<<40*5=200>>200 miles for work’ becomes ‘He bikes 202=<<202=40>>40 miles each day for work. So he bikes 405=<<405=200>>200 miles for work’). Table 4 comprehensively outlines each dataset, including its source and the size of the training, validation, and test samples.
Appendix E Removal Ratio Analysis on additional datasets.
@@
In the GSM8K, Strategy-QA, and Math-QA datasets, similar trends are observed as in Section 5.5.5. Notably, in Strategy-QA, the Removal Ratio converges more rapidly than in other datasets. This is attributed to the dataset’s characteristics, where the sentences within the rationale predominantly list simple information rather than forming complex, interrelated structures. Consequently, redundant sentences are rapidly removed, leading to a rapid decrease in further removal activities.
Appendix F Qualitative Analysis
Table 5 illustrates the generation outputs for each method on the GSM8K test set. Explicit-CoT produces a complete reasoning path for answer generation; however, its redundancy in generating ‘needs a 3 piece place setting’ leads to hallucinations within the sentence, resulting in the incorrect final answer generation. In contrast, ICoT-SI, which removes reasoning paths at the token-level, results in a generation that lacks sufficient reasoning capabilities. Conversely, VARR+ efficiently generates key information by producing ‘\$15.00’ in a single sentence, leading to more effective and efficient generation.
@@ Appendix F Qualitative Analysis
### 216

### v3

@@ Appendix A Additional Datasets Analysis
As shown in Section 3, the NLL tends to rise with increasing size of $\{r_{i}\}$ in random and back configurations, compared to the front configuration, as illustrated in Figure 6, this trend is consistent across both the GSM8K and StrategyQA datasets.
@@ Appendix B VARR Implementation
Appendix C Additional Experimental Details
For all experiments, we employ the AdamW optimizer Loshchilov and Hutter (2019), configured with a weight decay of 0.005. For the Mistral 7B model, we utilize an effective batch size of 12 with gradient accumulation set to 3, while the smaller Llama3.2 models use an effective batch size of 15. For Coconut (Hao et al., 2024), we set the max_latent_stage to 5 while maintaining all other hyperparameters as their default repository values unless otherwise mentioned. A constant learning rate of $5\times 10^{-6}$ is applied across all datasets, with bfloat16 precision. For Multiple-Choice and True/False tasks, complete sets of non-correct labels are employed to configure in-batch negatives to enhance the stability of the verbosity evaluation. To ensure a fair comparison, all baselines and methods are trained on a single A100 GPU with 80GB of memory for up to 5 epochs or 36 hours, whichever is reached first and experimented with single run evaluation (note that our setting is different from prompting/inference only setting). Regarding licensing, Mistral 7B is licensed under Apache License, Version 2.0, while Llama3.2 is governed by the Llama 3.2 Community License.
@@ Appendix D Details and Statistics of Datasets
GSM8K (Cobbe et al., 2021a) Arithmetic Cobbe et al. (2021b) Number 7000 473 1319
@@
StrategyQA (Geva et al., 2021) Commonsense Sileo (2024) T/F 1832 114 344
Table 4: Comprehensive statistics of the datasets used in our experiments are provided. GSM8K and MathQA are sourced directly from their original datasets, while the remaining datasets were obtained from Kim et al. (2023) and Sileo (2024) to access complete reasoning paths. GSM8K, StrategyQA, and CommonsenseQA are licensed under the MIT License, whereas MathQA and TriviaQA are distributed under the Apache License, Version 2.0.
For our experimental analysis, we carefully selected a diverse set of five datasets used in prior works (Deng et al., 2023, 2024a; Liu et al., 2024; Yu et al., 2024b; Yin et al., 2024). To ensure explicit sentence boundaries, all datasets were preprocessed using gpt-4o-mini (OpenAI, 2024) to establish clear sentence demarcation (e.g., ‘He bikes 20*2=<<20*2=40>>40 miles each day for work So he bikes 40*5=<<40*5=200>>200 miles for work’ becomes ‘He bikes 202=<<202=40>>40 miles each day for work. So he bikes 405=<<405=200>>200 miles for work’). Table 4 comprehensively outlines each dataset, including its source and the size of the training, validation, and test samples.
MQA G8K CQA TQA SQA
w/o reinit 48.70 45.18 87.82 83.72 76.74
(92.86) (115.91) (32.95) (38.27) (33.89)
w/ reinit 56.95 54.98 89.56 83.45 78.19
(91.04) (100.38) (33.55) (39.17) (35.12)
Table 5: Performance comparison before and after the reinitializing optimizer in every training epoch. Each row presents accuracy in the first line, with average generated tokens shown in parentheses in the second line.
Appendix E Reinitializing the Optimizer
We reinitialized the optimizer after each training epoch to stabilize training, inspired by Deng et al. (2024b). Our implementation uses the AdamW optimizer Loshchilov and Hutter (2019), where the first and second moments are gradually updated based on current gradients. Consequently, when VARR reduces rationales for certain data points between epochs, the training process could become unstable. As shown in table 6, the average 19.32% performance improvement achieved through optimizer reinitialization extends the findings of Deng et al. (2024b) beyond simple tasks (e.g., multiplication; synthesized dataset) to demonstrate effectiveness across diverse datasets with complex semantic and syntactic reasoning structures.
Appendix F Zero Warm-up Stage
warm-up 0.0 0.1 0.2 0.3 0.4
TQA 82.01 83.72 83.45 81.19 81.97
(40.11) (38.27) (47.10) (46.55) (47.97)
CQA 90.34 89.56 87.82 87.82 86.08
(32.81) (35.07) (39.38) (40.98) (43.61)
Table 6: Accuracy and average generated tokens across the various warm-up stages.
To validate whether the warm-up stage sensitivity could limit applicability, we conducted additional experiments extending warm-up stages to zero. As shown in Table, the zero warm-up stage shows comparable effectiveness and efficiency to the 0.1 warm-up stage, demonstrating the robustness of our method across different application scenarios. These findings suggest a zero-warm stage can be a good option for various application scenarios, but starting with a 0.1 warm-up stage can boost its stability.
Appendix G Removal Ratio Analysis on additional datasets.
@@ Appendix F Zero Warm-up Stage
In the GSM8K, StrategyQA, and MathQA datasets, similar trends are observed as in Section 5.5.4. Additionally, we observe that our method reduces fewer sentences in later training stages, demonstrating that verbosity preserves essential reasoning sentences even when larger $r(t)$ encourages substantial reduction. Notably, in StrategyQA, the Removal ratio converges more rapidly than in other datasets. This is attributed to the dataset’s characteristics, where the sentences within the rationale predominantly list simple information rather than forming complex, interrelated structures. Consequently, redundant sentences are rapidly removed, leading to a rapid decrease in further removal activities.
Appendix H Exploration on Complex Task
Exp-CoT VARR+
Acc 9.01 10.27
Gen Tok 300.17 252.00
Table 7: Performance comparison between Exp-CoT and VARR+ on MATH datasets.
To further validate that our method remains effective on complex tasks, we conducted additional experiments with the MATH (Hendrycks et al., 2021) dataset. Considering computational constraints, we selected 500 samples from the math subset of OpenMathInstruct-2 (Toshniwal et al., 2024) for the training set and evaluated on the official test set. As described in Table 8, our results on MATH further demonstrate the versatility and future applicability of VARR+ on more complex tasks.
Appendix I Training Time Analysis
Dataset GSM8K (7K) TQA (2K)
Training Time ($\uparrow$) 0.18 0.30
Token Efficiency ($\downarrow$) 0.81 0.61
Accuracy ($\uparrow$) 0.98 1.01
Table 8: Training time analysis on GSM8K and TQA. All scores are calculated as the value for Explicit-CoT divided by the value for VARR+ (i.e., Explicit-CoT/VARR+).
As VARR computes the likelihood of each sentence to ensure robust performance, training time inevitably increases as dataset size grows. We therefore quantitatively measure its efficiency on GSM8K and TQA relative to Explicit-CoT. Although there is a trade-off in training time, VARR+ improves token efficiency while preserving accuracy. Since training occurs only once while the improved efficiency benefits persist continuously afterward in test-time, we believe the additional training cost is justified.
Appendix J Qualitative Analysis
Table 9 illustrates the generation outputs for each method on the GSM8K test set. While Explicit-CoT generates a complete reasoning path, its redundant generation of ‘needs a 3 piece place setting’ introduces hallucinations into the reasoning process, ultimately leading to an incorrect answer. ICoT-SI’s token-level elimination of reasoning paths, on the other hand, results in generations that lack sufficient reasoning capacity. Additionally, Coconut’s predefined hidden-states fail to compress the complete reasoning path in its token-level compression approach. In contrast, VARR+ demonstrates efficient generation by concisely producing ‘\$15.00’ in a single sentence, achieving both effectiveness and efficiency in its output. Additional qualitative examples on the MathQA dataset are outlined in Table 10.
@@ Appendix I Training Time Analysis
#### 216
Coconut
<|start-latent|><|latent|><|latent|><|latent|><|end-latent|>
#### 4.00
@@
Question
The cost price of 20 articles is the same as the selling price of $x$ articles. if the profit is 25%, find out the value of x.
a) 13, b) 14, c) 15, d) 16, e) 17
Explicit-CoT
explanation : let the cost price 1 article = rs . 1
cost price of $x$ articles = rs . $x$
selling price of x articles = 125% of $x$ = rs . ( 125 / 100 $x$ ) = rs . ( 25 / 20 $x$ ) = rs . 5 / 4 $x$
= > $x$ = 20
cost price of 20 articles = rs . 20 selling price of 20 articles = rs . ( 25 / 4 $\times$ 20 ) = rs . ( 25 / 4 $\times$ 20 ) / 20 = rs . 1
= > $x$ = 20 answer : option a
#### a
ICoT-SI
of $x$ articles = $x$
s.p of $x$ articles = rs . 20
gain = rs . ( 20 - $x$ )
gain % = 25 % of $x$ = rs . ( 20 - $x$ ) / $x$$\times$ 100 = 25
= > 20 - $x$ = 5 $x$ = > 4 $x$ = 20 - $x$ = > 5 $x$ = 20 - $x$ = > 20 = 5 $x$ = > 200 = 5 $x^{2}$ = > 400 = 5 $x^{2}$ = > 400 / 5 = $x^{2}$ = > $x^{2}$ = 80 = > x = 10 or 10 . c
#### c
Coconut
<|start-latent|><|latent|><|latent|><|latent|><|end-latent|>
#### c
VARR+
explanation : let the cost price of $x$ articles = $x$
s.p of $x$ articles = 20
20 = 1.25 $x$ = > $x$ = 16 option d.
#### d
Ground Truth Answer
#### d
Table 10: Qualitative Example: Generation outputs on MathQA test set.